Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost

Direct materials3feet$11per foot$33
Direct labor?hours
?per hour
?

During March, the company purchased direct materials at a cost of $111,300, all of which were used in the production of 3,200 units of product. In addition, 4,900 direct labor-hours were worked on the product during the month. The cost of this labor time was $95,550. The following variances have been computed for the month:

 

Materials quantity variance$4,400U
Labor spending variance$450F
Labor efficiency variance$2,000U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

 

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1

ANSWER:

1) Material quantity variance = (Standard quantity-actual quantity)Standard price

-4400 = (3200*3*11-11*X)

-4400 = 105600-11*X

11*X = 110000

X(actual quantity used) = 110000/11

= 10000 pounds

Actual cost per foot = 111300/10000

= 11.13 per foot

2) Material price variance = (11*10000-111300)

= 1300 U

Material spending variance = 3200*33-111300

= 5700 U

3) Labor rate variance = 2000+450

= 2450 F

Labor rate variance = (Standard rate-actual rate)actual hours

2450 = (4900*X-95550)

-4900*X = -98000

X(Standard rate) = 20 per hour

4) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-2000 = (X*20-4900*20)

-2000 = 20*X-98000

-20*X = -96000

X(standard hour) = 4800 Hour

5) Standard hour per unit = 4800/3200

= 1.5 per hour

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************